OpenStudy (anonymous):

plzzz help me Q6 The circuit elements have the following values: Vin=18V, R1=1.5Ω, R2=3Ω, R3=4Ω, A=1.5 and B=0.5. What is the value of the current i1 (in Amps)? What is the value of the output voltage vOUT (in Volts)? We wish to create a Thevenin equivalent model of the above shown circuit as seen from its Output Port. What is the value of the Thevenin equivalent voltage as seen from the Output Port (in Volts)? What is the value of the Thevenin equivalent resistance (in Ohms) as seen from the Output Port?

OpenStudy (anonymous):

i1=1.8 vOUT=1.8 Vth=1.8 Rth=3.6

OpenStudy (anonymous):

neelima plz help me too, plz send the formula.

OpenStudy (anonymous):

i1=vOUT/(R3-(B*R1*R3)) vOUT=Vin(R3-(B*R1*R3))/(R3-(B*R1*R3)+R1+R2+(A*R2)) Vth=vOUT Rth=Vth/Isc where Isc=Vin*(1-(B*R1))/((A*R2)+R1+R2)

OpenStudy (anonymous):

Vin=45V, R1=1.5Ω, R2=2Ω, R3=6Ω, A=2 and B=0.5.

OpenStudy (anonymous):

neelima can u calclulate for me?

OpenStudy (anonymous):

i1=5 vOUT=7.5 Vth=7.5 Rth=5

OpenStudy (anonymous):

Thanku neelima

OpenStudy (anonymous):

Q1 R1=50Ω, R2=50Ω, R3=100Ω, V1=4V, I1=0.0001A and I2=0.001A (a) What is the value of voltage vA in Volts? (b) What is the value of voltage vB in Volts? Now let us determine if the answers you came up with satisfy the laws of physics. What is the power (in Watts) dissipated in resistor R1? What is the power (in Watts) dissipated in resistor R2? What is the power (in Watts) dissipated in resistor R3? ? ?

OpenStudy (anonymous):

Neelima Plzzz help me again

OpenStudy (anonymous):

vA = (R2*V1 + R3*V1 + I1*R1*R3 - I2*R1*R3)/(R1 + R2 + R3) vB = (R3*(V1 + I1*R1 + I1*R2 - I2*R1 - I2*R2))/(R1 + R2 + R3) Power for R1 = (vA-V1)^2/R1 Power for R2 = (vA-vB)^2/R2 Power for R3 = vB^2/R3 Power for V1 = V1 * (V1-vA)/R1 Power for I1 = V1 * I1 Power for I2 = V1* (-I2)

OpenStudy (anonymous):

Q2 VS=5.0V, VOH=4.5V, VIH=4.0V, VIL=1.5V, VOL=1.0V. What is the low noise margin (in Volts)? What is the high noise margin (in Volts)? What is the width of the forbidden region (in Volts)? Suppose that the threshold voltage for the MOSFET is VT=2.0V and RON=16000.0Ω. What is the minimum value of the pullup resistor RPuI (in Ohms) for which this inverter can obey the required static discipline? Now, consider the NAND gate of this family. What is the minimum value of the pullup resistor RPuA (in Ohms) for which this inverter can obey the required static discipline? How about the NOR gate of this family. What is the minimum value of the pullup resistor RPuO (in Ohms) for which this inverter can obey the required static discipline? Assume that we implemented this family with the minimum pullup resistors that you have already calculated. What is the maximum power (in Watts) consumed by the inverter? What is the maximum power (in Watts) consumed by the NAND? What is the maximum power (in Watts) consumed by the NOR?

OpenStudy (anonymous):

@apurvdixit713 plzzz values of resistences Rpul etc

OpenStudy (anonymous):

low noise margin=VIL-VOL=0.5 HIGH NOISE MARGIN=VOH-VIH=0.5 forbidden region=VIH-VIL=2.5 RpuI=4*RON=24000 RpuA=8*RON=48000 RpuO=4*RON=24000

OpenStudy (anonymous):

power for inverter=? power for NAND=? power for NOR=?

OpenStudy (anonymous):

if VT=2.0V and RON=16000.0Ω.

OpenStudy (anonymous):

VT=2.0V and RON=16000.0Ω. power for inverter=? power for NAND=? power for NOR=?

OpenStudy (anonymous):

@apurvdixit713 What is the maximum power (in Watts) consumed by the inverter? Answer : Power = (Vs^2)/ (Rpull + Ron total) What is the maximum power (in Watts) consumed by the NAND? Answer : Power = (Vs^2)/ (Rpull + 2*Ron total) What is the maximum power (in Watts) consumed by the NOR? Answer : Power = (Vs^2)/ (Rpull + Ron||Ron)

OpenStudy (anonymous):

Ron =v Ron total . . for you its 16000

OpenStudy (anonymous):

Ron = Ron total . . . 16000

OpenStudy (anonymous):

@apurvdixit713 0.000625 0.0003125 0.00078125

OpenStudy (anonymous):

@apurvdixit713 got ittt . .

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

u have checked it??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

resistances value correct??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

power for inverter=0.000833 power for NAND=0.000416 power for NOR=0.000925

OpenStudy (anonymous):

how many checks u have??

OpenStudy (anonymous):

@apurvdixit713 how many checks do u have??

OpenStudy (anonymous):

last one chance

OpenStudy (anonymous):

@apurvdixit713 wait just 5 mins

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

it is correct formula or not What is the maximum power (in Watts) consumed by the inverter? Answer : Power = (Vs^2)/ (Rpull + Ron total) What is the maximum power (in Watts) consumed by the NAND? Answer : Power = (Vs^2)/ (Rpull + 2*Ron total) What is the maximum power (in Watts) consumed by the NOR? Answer : Power = (Vs^2)/ (Rpull + Ron||Ron)

OpenStudy (anonymous):

@apurvdixit713 i m back . . wait

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

these r corrct 0.000625 0.0003125 0.00078125

OpenStudy (anonymous):

after putting click save and then check button

OpenStudy (anonymous):

sure it is wright

OpenStudy (anonymous):

@apurvdixit713 . . now Q2 done??

OpenStudy (anonymous):

bcz this is my last chance

OpenStudy (anonymous):

just do it .. after putting all the values . . save then check

OpenStudy (anonymous):

again its show wrong.

OpenStudy (anonymous):

have you clicked save . . it will show the last answer . .then click chek

OpenStudy (anonymous):

for RON=16000....... RpuI=64000 RpuA=128000 RpuO=64000 power for Inverter=0.0003125 power for NAND=0.000156 power for NOR=0.000347 check this once...